# REvolve: Reward Evolution with Large Language Models using Human Feedback

### v2

@@ 3 Reward Evolution
As the fitness function $F$ is based on implicit human understanding, it cannot be explicitly formulated and is, therefore, not amenable to gradient-based optimization. Hence, we solve the RDP using a meta-heuristic approach through genetic programming (GP) (Langdon & Poli, 2013). Concretely, REvolve uses GP to output and iteratively refine a population of reward functions (individuals) by applying genetic operations such as mutation, crossover, and migration. In each iteration, we obtain new individuals, which are added to the population based on their fitness. This fitness is assessed through human evaluations of the policies trained with these reward functions – thus guiding the evolutionary search.
Evolutionary algorithms, in general, are considered to be weak methods that perform blind searches and do not exploit domain-specific knowledge (De Jong, 2017). The initialization and perturbations are random, leading to scalability issues (O’Neill et al., 2010). In contrast, LLMs can leverage their commonsense priors to suggest changes that are intelligent and better aligned with the feedback (Gugerty & Olson, 1986). Therefore, we combine the evolutionary search of REvolve with LLMs by using them as reward designers $G$ in the RDP – thus avoiding blindly searching in the solution space.
@@ 4 Related Work
Reward Design with LLMs. Recent works have explored LLMs for complex search and black-box optimization problems like hyperparameter optimization (Yang et al., 2024; Zhang et al., 2023; Liu et al., 2024), reward design for RL (Kwon et al., 2023; Xie et al., 2024; Ma et al., 2024a), or robotic control (Yu et al., 2023). Kwon et al. (2023) use LLMs to directly output binary rewards by inferring user intentions in negotiation games, but this requires numerous queries to LLMs and limits the rewards to sparse binary values. Conversely, Language to Rewards (L2R) (Yu et al., 2023) and Text2Reward (T2R) (Xie et al., 2024) utilize LLMs to create dense and interpretable reward functions, improving them with human feedback. L2R, however, restricts its adaptability by employing manually crafted reward templates tailored to specific robot morphologies. In contrast, T2R outputs versatile reward functions in Python code. It generates a single reward function per iteration, checks for improvement from the previous one, and repeats this process, hence performing a greedy search over iterations.
Eureka (Ma et al., 2024a) further extends T2R by generating multiple reward functions and greedily iterating with the best reward function and discarding the rest. This means that by definition (Lehman et al., 2024), Eureka does not qualify as an EA since it lacks crucial features. Namely, maintaining and evolving a population of candidates and employing genetic operators like mutation, crossover, and selection. Eureka only mutates the best candidate and discards the rest. In contrast, REvolve is truly evolutionary, i.e., population-based and using the full range of genetic operators. By preserving genetic diversity, REvolve avoids the pitfalls of premature convergence without incurring additional costs (since both frameworks generate the same number of candidates per generation). Empirically, we show that utilizing the full range of evolutionary operators yields reward functions that lead to better performing agents. Consequently, REvolve significantly outperforms greedy search methods like Eureka and T2R.
@@ 5.2 Training Setup
For the evolutionary search, we set the number of generations to $N=5$ and individuals per generation to $K=16$. This mimics the training setup of (Ma et al., 2024a). The number of sub-populations was set to $I=13$, and the mutation probability was set to $p_{m}=0.5$.
@@
In addition to REvolve (with human feedback), we also include a version in the experimental evaluation that we dub REvolve Auto. Here, the human feedback is replaced with automatically generated feedback comprised of statistics on the reward components tracked during RL training. Moreover, the fitness score, evaluating the individual agents, is computed using a manually designed fitness function. This fitness function measures an agent’s successful completion of the task at hand. For instance, the fitness score for the humanoid locomotion task encodes that the agent should not fall over and that it should attain a high average velocity. We formally describe the fitness scores for the three different environments in Appendix B.2. This automated feedback mechanism is adopted from (Ma et al., 2024a).
@@ 5.4 Experimental Questions & Results
We also note that in the Humanoid Locomotion and Adroit Hand tasks REvolve Auto (i.e., with automated feedback) performs on par with Eureka (with human feedback), despite the former’s evolutionary setting. This could be attributed to the limited and sparse nature of automated feedback (velocity for humanoid, steps to succeed for adroit hand) in contrast to the more detailed human feedback in Eureka. In contrast, automated feedback in Autonomous Driving is dense and incorporates multiple factors like collision, speed, and distance (Appendix B.2) – leading to bigger improvements during the evolutionary search. This also highlights that the combination of evolutionary search and human feedback is crucial for optimizing reward functions.
@@ 5 Expert Designed 1415
We further compared REvolve and Eureka on human evaluations. To this end, we generated roll-outs from the best performing policies ($\pi^{\ast}$) from REvolve and Eureka. We paired them against each other and used the user feedback interface (Figure 1) for collecting preference data. This data was subsequently passed through an Elo rating system to rank each behavior, cf. Appendix B.1. As shown in Table 2, REvolve policies again outperform Eureka policies across all three tasks, and it is only the human expert driver beating REvolve in Autonomous Driving.
We have submitted example rollout videos as part of the supplementary material.
@@ 5.5 Ablation Studies
Figure 4: Comparison of REvolve’s performance on Autonomous Driving using mutation only, crossover only, and both combined.
@@ 6 Conclusion & Future Work
During human feedback experiments, we engaged 10 human evaluators per generation, each assessing 20 pairs of videos. Participants were recruited on a volunteer basis and the preference learning task was clearly explained to them. We note that no personal data was collected, and participants were compensated fairly per institutional guidelines. Our study adhered to ethical standards, followed institutional guidelines, and did not involve any potential risks related to privacy, security, or discrimination.
@@
In the Appendix, we provide extensive details on our experimental protocol. Specific details can be found in the respective sub-appendices as detailed below. The reinforcement learning framework, including the algorithm, inputs, outputs, and hyperparameters, is outlined in Appendix A. The fitness functions are thoroughly discussed in Appendix B. The expert-designed rewards are explained in Appendix C. Detailed prompt examples for GPT-4 can be found in Appendix E. Finally, we list the best reward functions generated from REvolve in Appendix G. We will open-source our code shortly.
@@ Appendix D REvolve Cost Analysis & Comparisons
Data requirement: RLHF >>much-greater-than$>$ > REvolve: RLHF requires massive amounts of curated data to train a reward model. Note that the input to a reward model is a video (rollout of a policy), and the output is a preference score, requiring a heavy neural model. As a reference, the common video classification model I3D trains on the Kinetics dataset, which has 650,000 video clips (Carreira & Zisserman, 2017). IRL >>much-greater-than$>$ > REvolve: Similarly, Inverse Reinforcement Learning requires human demonstrations and interventions. In contrast, REvolve uses 10 human evaluators per generation, each assessing 20 data samples (video pairs) – hence, for 5 generations $10\times 20\times 5=1000$ curated samples only. Compare this to the Lyft self-driving dataset (Houston et al., 2021) used for motion prediction which has more than 1000 hours of data collected over 4 months by 20 cars.
Storage Cost: REvolve $>$ Eureka. The slightly higher cost is due to maintaining a population of best candidates rather than one candidate like Eureka. However, compared to the LLM memory cost (e.g., the biggest open-source LLM Llama 3.1 takes up 810 GB of memory), the storage cost is negligible – for every generation, the storage increases by  2.5GB (which includes stored policy weights and other metadata) – 12.5 GB after 5 generations. Also, REvolve does not require expert feedback.

### v3

@@ 3 Reward Evolution
As the fitness function $F$ is based on implicit human understanding, it cannot be explicitly formulated and is, therefore, not amenable to gradient-based optimization. Hence, we solve the RDP using a meta-heuristic approach through genetic programming (GP) (Langdon & Poli, 2013). Concretely, REvolve uses GP to iteratively refine a population of reward functions (individuals) by applying genetic operations such as mutation, crossover, and migration. In each iteration, new individuals are obtained and added to the population based on their fitness. This fitness is assessed through human evaluations of the policies trained with these reward functions – thus guiding the evolutionary search.
Evolutionary algorithms, in general, are considered to be weak methods that perform blind searches and do not exploit domain-specific knowledge (De Jong, 2017). The initialization and perturbations are random, leading to scalability issues (O’Neill et al., 2010). In contrast, LLMs can leverage their commonsense priors to suggest changes that are intelligent and better aligned with the feedback (Gugerty & Olson, 1986). Therefore, we combine the evolutionary search of REvolve with LLMs by using them as reward designers $G$ in the RDP.
@@ 4 Related Work
Reward Design with LLMs. Recent works have explored LLMs for complex search and black-box optimization problems like hyperparameter optimization (Yang et al., 2024; Zhang et al., 2023; Liu et al., 2024), reward design for RL (Kwon et al., 2023; Xie et al., 2024; Ma et al., 2024a), or robotic control (Yu et al., 2023). Kwon et al. (2023) use LLMs to directly output binary rewards by inferring user intentions in negotiation games, but this requires numerous queries to LLMs and limits the rewards to sparse binary values. Conversely, Language to Rewards (L2R) (Yu et al., 2023) and Text2Reward (T2R) (Xie et al., 2024) utilize LLMs to create dense and interpretable reward functions, improving them with human feedback. L2R, however, restricts its adaptability by employing manually crafted reward templates tailored to specific robot morphologies. In contrast, T2R outputs versatile reward functions in Python code. It generates a single reward function per iteration, checks for improvement from the previous one, and repeats this process, hence performing a greedy search.
Eureka (Ma et al., 2024a) further extends T2R by generating multiple reward functions and greedily iterating with the best reward function and discarding the rest. This means that by definition (De Jong, 2017), Eureka does not qualify as an EA since it lacks crucial features. Namely, maintaining and evolving a population of candidates and employing genetic operators like mutation, crossover, and selection. Eureka only mutates the best candidate and discards the rest. In contrast, REvolve is truly evolutionary, i.e., population-based and using the full range of genetic operators. By preserving genetic diversity, REvolve avoids the pitfalls of premature convergence without incurring additional costs (since both frameworks generate the same number of candidates per generation). Empirically, we show that utilizing the full range of evolutionary operators yields reward functions that lead to better performing agents. Consequently, REvolve significantly outperforms greedy search methods like Eureka and T2R.
@@ 5.2 Training Setup
For the evolutionary search, we set the number of generations to $N=7$ and individuals per generation to $K=16$. This mimics the training setup of (Ma et al., 2024a). The number of sub-populations was set to $I=13$, and the mutation probability was set to $p_{m}=0.5$.
@@
In addition to REvolve (with human feedback), we also include a version in the experimental evaluation that we dub REvolve Auto. Here, the human feedback is replaced with automatically generated feedback comprising statistics on the reward components tracked during RL training. The fitness score for each agent is now computed using a manually designed fitness function, which measures an agent’s successful completion of the task. For instance, the fitness score for the humanoid locomotion task encodes that the agent should not fall over and that it should attain a high average velocity. We formally describe the fitness scores for the three different environments in Appendix B.2. This automated feedback mechanism is adopted from (Ma et al., 2024a).
@@ 5.4 Experimental Questions & Results
We also note that in the Humanoid Locomotion and Adroit Hand tasks REvolve Auto (i.e., with automated feedback) performs on par with Eureka (with human feedback), despite the former’s evolutionary setting. This could be attributed to the limited and sparse nature of automated feedback (velocity for humanoid, steps to succeed for adroit hand) in contrast to the more detailed human feedback in Eureka. In contrast, automated feedback in Autonomous Driving is dense and incorporates multiple factors like collision, speed, and distance (Appendix B.2) – leading to bigger improvements during the evolutionary search. This also highlights that the combination of evolutionary search and human feedback is crucial for optimizing reward functions. It can be observed that the improvements gradually diminish with each generation and stabilize, converging by generation 5.
@@ 5 Expert Designed 1415
We further compared REvolve and Eureka on human evaluations. To this end, we generated roll-outs from the best performing policies ($\pi^{\ast}$) from REvolve and Eureka. We paired them against each other and used the user feedback interface (Figure 1) for collecting preference data. This data was subsequently passed through an Elo rating system to rank each behavior, cf. Appendix B.1. As shown in Table 2, REvolve policies again outperform Eureka policies across all three tasks, and it is only the human expert driver beating REvolve in Autonomous Driving. Note, we used a fresh batch of evaluators (disjoint from training) for this stage. We have submitted example rollout videos as part of the supplementary material.
@@ 5.5 Ablation Studies
Figure 4: Comparison of REvolve’s performance on Autonomous Driving using mutation only, crossover only, and both combined, over 7 generations.
@@ 6 Conclusion & Future Work
During human feedback experiments, we engaged 10 human evaluators per generation, each assessing 20 video pairs. Participants were recruited on a volunteer basis and the preference learning task was clearly explained to them. No personal data was collected, and participants were compensated fairly per institutional guidelines. Our study adhered to ethical standards, followed institutional guidelines, and did not involve any potential risks related to privacy, security, or discrimination.
@@
In the Appendix, we provide extensive details on our experimental protocol. Specific details can be found in the respective sub-appendices as detailed below. The reinforcement learning framework, including the algorithm, inputs, outputs, and hyperparameters, is outlined in Appendix A. The fitness functions are thoroughly discussed in Appendix B. The expert-designed rewards are explained in Appendix C. Detailed prompt examples for GPT-4 can be found in Appendix E. Finally, we list the best reward functions generated from REvolve in Appendix G. Our code is open-sourced at https://github.com/RishiHazra/Revolve/tree/main.
Acknowledgments
The computations and data handling for this work were enabled by the Berzelius resource at the National Supercomputer Centre, provided by the Knut and Alice Wallenberg Foundation, under projects “Berzelius-2024-178” and “Berzelius-2023- 348”. This research was supported by the Wallenberg AI, Autonomous Systems, and Software Program (WASP), also funded by the Knut and Alice Wallenberg Foundation. Additionally, this work is part of the NEST Project titled “Multi-dimensional Alignment and Integration of Physical and Virtual Worlds” (__main__). Finally, we thank Luc De Raedt for his valuable feedback in refining the manuscript.
@@ Appendix D REvolve Cost Analysis & Comparisons
Data requirement: RLHF >>much-greater-than$>$ > REvolve: RLHF requires massive amounts of curated data to train a reward model. Note that the input to a reward model is a video (rollout of a policy), and the output is a preference score, requiring a heavy neural model. As a reference, the common video classification model I3D trains on the Kinetics dataset, which has 650,000 video clips (Carreira & Zisserman, 2017). IRL >>much-greater-than$>$ > REvolve: Similarly, Inverse Reinforcement Learning requires human demonstrations and interventions. In contrast, REvolve uses 10 human evaluators per generation, each assessing 20 data samples (video pairs) – hence, for 7 generations $10\times 20\times 7=1400$ curated samples only. Compare this to the Lyft self-driving dataset (Houston et al., 2021) used for motion prediction which has more than 1000 hours of data collected over 4 months by 20 cars.
Storage Cost: REvolve $>$ Eureka. The slightly higher cost is due to maintaining a population of best candidates rather than one candidate like Eureka. However, compared to the LLM memory cost (e.g., the biggest open-source LLM Llama 3.1 takes up 810 GB of memory), the storage cost is negligible – for every generation, the storage increases by  2.5GB (which includes stored policy weights and other metadata) – 17.5 GB after 7 generations. Also, REvolve does not require expert feedback.